OpenStudy (anonymous):

The arithmetic mean of two numbers is 18. Their geometric mean is 16. Find the sum of the squares of the two numbers.

OpenStudy (anonymous):

arthemtic mean is \[(x+y)\div2\]=18 and geometric mean is xy=16 solve these equations for x and y

OpenStudy (anonymous):

what does it mean by the sum of the two squares?

OpenStudy (anonymous):

the sum of the two squares basically means that the sum of the two numbers after they have been squared.

OpenStudy (anonymous):

so far I have \[a+b=36\] & \[ab=256\] my teacher said that the geometric mean is\[\sqrt{ab}\]

OpenStudy (anonymous):

hey sry its \[\sqrt{xy}\]

OpenStudy (anonymous):

they mean \[x^{2}+y^{2}\]

OpenStudy (anonymous):

then u have a+b square it

OpenStudy (anonymous):

it will be \[a^{2}+b^{2}+2ab\]

OpenStudy (anonymous):

use ur values to obtain solution

OpenStudy (anonymous):

\[\left( 36 \right) \left( a^2+512+b^2 \right)\] is this the right equation?

OpenStudy (anonymous):

[36^{2}=a^{2}+b^{2}+512\]

OpenStudy (anonymous):

therefore, \[a^2+b^2=784\]

OpenStudy (anonymous):

haa....thats all hope u understood the solutionn

OpenStudy (anonymous):

Yes, thank you so much!